# Unlocking the Potential of Model Calibration in Federated Learning

### v2

@@ 1 Introduction
Federated learning (FL) has rapidly gained traction as a prominent distributed machine learning approach, enabling collaborative model training across a network of clients by periodically aggregating their local models at a central server [Konecný et al., 2016; McMahan et al., 2017]. Its significance extends to various critical applications, including medical diagnostics, self-driving cars, and multilingual systems, where data privacy and decentralized training are paramount. FL has been widely studied with a focus on improving local model updates Reddi et al. [2021]; Sahu et al. [2018], refining global aggregation methods Ji et al. [2019]; Wang et al. [2020a], and enhancing communication efficiency Sattler et al. [2019]; Diao et al. [2020]. Most of these works in FL consider accuracy as the main performance metric.
@@
Motivation. However, beyond accuracy, in various decision-making scenarios where an incorrect prediction may result in high risk (e.g., medical applications or autonomous driving), it is also crucial for the users to determine whether to rely on the FL model’s prediction or not for each decision. In particular, users should rely on the neural network’s decision only when the prediction is likely to be correct. Otherwise, they may need to consider alternatives such as human decision. To achieve this, the trained FL model should have a reliable confidence in each of its predictions, meaning that the confidence of the neural network matches well with its actual accuracy. In centralized settings, recent studies including Guo et al. [2017] have uncovered that neural networks are often miscalibrated, indicating that the prediction confidence of the model does not accurately reflect the probability of correctness. Handling this miscalibration issue is even more important in many FL use-cases, where an overconfident global model could lead to misinformed decisions with potentially severe consequences for each client.
@@
In centralized training settings, research generally follows two paths to address this miscalibration issue. First, train-time calibration methods Hebbalaguppe et al. [2022b]; Liang et al. [2020]; Liu et al. [2021]; Mukhoti et al. [2020]; Kumar et al. [2018]; Müller et al. [2019]; Lin et al. [2017] incorporate explicit regularizers during the training process to adjust neural networks, effectively scaling back over/under-confident predictions. On the other hand, post-hoc calibration Guo et al. [2017]; Hendrycks & Gimpel [2016]; Hinton et al. [2015]; Kull et al. [2017] transforms the network’s output vector to align the confidence of the predicted label with the actual likelihood of that label for the sample. Post-hoc calibration methods are applied to the already trained model to improve calibration using an additional holdout dataset. However, an appropriate holdout dataset may not be available in many privacy-sensitive applications, making post-hoc calibration strategies often impractical for FL.
@@
Figure 1: Reliability diagrams and calibration errors for centralized training and non-IID FL (using FedAvg) trained with various calibration methods. Our method ensures well-calibrated FL, evidenced by a notably smaller calibration error and a smaller gap (red region) between confidence and accuracy.
@@
Goals and observations. In this paper, we aim to incorporate model calibration into the FL training process, a problem that has been largely underexplored in existing research. This setting introduces new challenges to improving model calibration, as the heterogeneous data distributions across different clients must be carefully considered. We first observe whether existing train-time calibration methods applied during client-side local training can effectively address calibration needs in FL. To gain insights, in Fig. 1, we report accuracy, confidence, and the expected calibration error (ECE) of different methods in different settings. Figs. 1(a) to  1(b) illustrate the capability of train-time calibration methods to narrow the gap between model confidence and accuracy in centralized learning. A comparison between Figs. 1(a) and  1(c) reveals that FL experiences more significant model miscalibration than centralized learning. This disparity is likely due to data heterogeneity across distributed FL clients, causing each client’s local data to have a different impact on calibration performance. When applying auxiliary-based calibration methods, such as DCA Liang et al. [2020] and MDCA Hebbalaguppe et al. [2022b] – which add a penalty term to the original classification loss – to FL’s local training (Fig. 1(d)), a better ECE is achieved compared to FL without calibration (Fig. 1(c)). However, we observe that the model in Fig. 1(d) is still not well-calibrated by neglecting global calibration needs in heterogeneous FL settings, resulting in overconfident predictions. We thus pose the following question:
@@
Contributions. In this paper, we propose Non-Uniform Calibration for Federated Learning (NUCFL), a versatile and generic model calibration framework for FL that integrates seamlessly with any existing FL algorithms, such as FedProx Sahu et al. [2018], Scaffold Karimireddy et al. [2019], FedDyn Acar et al. [2021], and FedNova Wang et al. [2020b]. Designed to work with any train-time calibration auxiliary loss, NUCFL is a straightforward yet effective solution that dynamically adjusts calibration penalties based on the relationship between global and local distribution. Specifically, our intuition is that if each client’s local model closely resembles the global model, it is likely that the local model and data represent the global calibration needs well. Therefore, our idea is to choose a large penalty term to improve model calibration for the clients that have a high similarity, while imposing a small penalty for other clients to prioritize accuracy. This non-uniform calibration approach considering data heterogeneity across FL clients significantly minimizes calibration errors without accuracy degradation, as seen in Figure 1(e). Importantly, the flexibility of NUCFL allows for continuous improvement as new FL algorithms or calibration losses are developed. Our key contributions are summarized as follows:
  • Our work is among the early efforts to conduct a systematic study on model calibration in FL. We study the question: How does model calibration impact the behavior of federated optimization methods?
@@
  • We conduct extensive experiments demonstrating that NUCFL is highly flexible, seamlessly integrating with 5 different FL algorithms (e.g., FedProx, Scaffold, FedDyn) and various auxiliary-based train-time calibration methods to improve both calibration performance and accuracy.
@@ 2 Related Work
Federated learning. FedAvg McMahan et al. [2017] is the pioneering FL algorithm, with numerous adaptations proposed to mitigate the accuracy drop experienced with non-IID data. Much of the research has focused on designing better global aggregation methods to enhance model convergence and performance. Some researchers He et al. [2020]; Lin et al. [2020] replaced weight averaging with model ensemble and distillation, others Yurochkin et al. [2019]; Wang et al. [2020a] matched local model weights before averaging, and another group Ji et al. [2019] introduced an attention mechanism to optimize global aggregation. Another significant area has concentrated on optimizing local training protocols to optimize computation and data usage on clients’ devices. Some researchers Sahu et al. [2018]; Acar et al. [2021] employed regularization toward the global model, while [Karimireddy et al., 2019] utilized server statistics to correct local gradients. Additionally, various studies have addressed challenges specific to FL, such as ensuring fairness across all participants Li et al. [2020], enhancing communication efficiency to minimize bandwidth usage Ko et al. [2023], and addressing privacy concerns inherent in distributed computations Truong et al. [2020]. However, the main focus of existing FL research is on improving accuracy. The mismatch between confidence and actual accuracy in these works leads to overconfident FL models, making FL inapplicable in high-risk decision-making scenarios. Our work bridges this gap by investigating an overlooked aspect of federated learning: model calibration.
@@
Model calibration. Model calibration techniques can be categorized into post-hoc calibration and train-time calibration methods. Post-hoc methods adjust the model using a hold-out dataset after training. Common examples include Temperature Scaling Guo et al. [2017], which adjusts logits by a scalar learned from a hold-out set before softmax, and Dirichlet Calibration Kull et al. [2017], which modifies log-transformed class probabilities through an additional network layer. However, in FL applications, an additional validation set is often not available due to privacy issues, making post-hoc calibration methods inapplicable. On the other hand, train-time calibration integrates calibration directly within the training process. Muller et al. Müller et al. [2019] enhance calibration using label smoothing on soft targets, while Lin et al. Lin et al. [2017] use focal loss to minimize the KL divergence between predicted and target distributions. Recently, augmenting standard cross-entropy loss with additional penalty loss terms has become a popular train-time calibration method Liang et al. [2020]; Hebbalaguppe et al. [2022b]; Kumar et al. [2018]. This auxiliary loss penalizes the model when there is a reduction in cross-entropy loss without a corresponding increase in accuracy, often indicative of overfitting. However, as seen in Fig. 1, a direct application of these methods still faces limitations, as they are not able to meet the calibration needs considering data heterogeneity across FL clients. Our work fills this gap by taking a non-uniform calibration approach considering the characteristics of FL.
Calibration and Federated Learning
@@
Peng et al. [2024] proposed FedCal, a calibration scaler aggregated solely from local clients, and introduced model calibration for FL. However, FedCal does not consider the interaction between global and local calibration needs, which can lead to a calibration bias towards local heterogeneity. Our model calibration framework addresses this issue by considering the interaction between client and server, thereby better accommodating broader global calibration requirements. Furthermore, while FedCal has shown effectiveness only with FedAvg, our framework aims to provide robust calibration across various FL configurations, enhancing flexibility beyond just FedAvg. It is also worth mentioning that Zhang et al. [2022]; Luo et al. [2021] also use the term “calibration” in the FL setup. However, we stress that the context of calibration in their work is different from the one in the model calibration literature that considers confidence of the prediction. Zhang et al. [2022]; Luo et al. [2021] focuses on calibrating the classifier to improve the accuracy itself, which can be categorized into conventional FL works that do not consider the error between model confidence and accuracy.
@@ 3.1 Federated Learning
In Federated Learning (FL), the training data is gathered from a set of clients $M$ and each client $m$ possesses a training set $\mathcal{D}_{m}=\{x_{i},y_{i}\}^{|\mathcal{D}_{m}|}_{i=1}$, where $x$ represents the input data and $y$ denotes the corresponding true label. We consider the following standard optimization formulation of federated training and seek to find model parameters $w$ that solve the problem:
@@
where $w$ represents the model parameter, $\mathcal{D}=\cup_{m}\mathcal{D}_{m}$ denotes the aggregated training set from all clients, and $\mathcal{L}_{m}$ measures the average loss of a model on the training data of the $m$-th client. The objective is to find a model that fits all clients’ data well on (weighted) average.
FL methods involve parallel local training at clients and global aggregation at a server over multiple communication rounds to address the aforementioned problem. Most federated optimization methods fit within a general framework outlined by [Reddi et al., 2020], as shown in Algorithm 1. At round $t$, the server sends its previous global model ${w}^{(t-1)}$ to all clients as initialization. Each client $m$ then performs $E$ epochs of local training using ClientOPT depending on the FL algorithm, and subsequently produces a local model ${w}_{m}^{(t,E)}$. Then each client communicates the difference $\delta^{(t)}_{m}$ between their learned local model and the server model, denoted as $\delta^{(t)}_{m}={w}^{(t-1)}-{w}^{(t,E)}_{m}$. The server computes a weighted average of the client updates, denoted as $\delta^{(t)}$, and then updates the global model ${w}^{(t)}$ using ServerOPT, which also depends on FL designs. For example, FedAvg employs standard stochastic gradient descent updates for ClientOPT, and its ServerOPT is formulated as $w^{(t)}=w^{(t-1)}-\delta^{(t)}$.
@@ 3.2 Model Calibration
In supervised multi-class classification, the input $x\in\mathcal{X}$ and label $y\in\mathcal{Y}=\{1,...,K\}$ are random variables following a ground truth joint distribution $\pi(x,y)=\pi(y|x)\pi(x)$. Let $s$ be the confidence score vector in $\mathbb{R}^{K}$, with $s[y]=f(x)$ representing the confidence that the neural network $f(\cdot)$ predicts for a class $y$ given input $x$. The predicted class $\hat{y}$ is computed as: $\hat{y}=\arg\max_{y}s[y]$, with the corresponding confidence $\hat{s}=\max_{y}s[y]$. A model is considered perfectly calibrated Guo et al. [2017] if:
@@
The primary goal of model calibration, whether via train-time or post-hoc methods, is to minimize discrepancy (3) to ensure the model’s confidence accurately reflects its prediction accuracy.
@@ 4.1 Bridging Federated Learning and Model Calibration
where $w_{m}$ is the local model, $t$ the current global round, $e$ the local iteration index, and $\ell(\cdot)$ the loss function (e.g., cross-entropy) applied to a data instance. One can also modify the loss in (4), i.e., by adding a proximal term as in FedProx Sahu et al. [2018]. Recent studies on calibration Liang et al. [2020]; Hebbalaguppe et al. [2022b] suggest that incorporating an auxiliary loss, which captures the disparity between accuracy and confidence (as in equation (3)), and penalizing the model accordingly, leads to superior performance compared to direct loss type adjustments Müller et al. [2019]; Brier [1950]; Lin et al. [2017]. These auxiliary loss methods maintain the original classification loss and seamlessly add a penalty, minimizing interference with FL performance, especially since some FL algorithms are specifically designed around their classification loss strategies. This makes the auxiliary loss methods easily applicable and effective across various FL settings.
@@
where $\ell_{cal}(\cdot)$ is the auxiliary calibration loss function, and $\beta_{m}$ is a scalar weight for each client $m$. In the centralized calibration case, this scalar weight is manually selected to refine the auxiliary calibration loss. However, in our design, we automatically determine $\beta_{m}$ based on the relationship between the server and clients in a FL setup, taking into account the heterogeneity as detailed in Section 4.2. The auxiliary calibration loss $\ell_{cal}$ can be derived from various calibration designs. For instance, [Liang et al., 2020] directly adds the difference between confidence and accuracy (DCA) as an auxiliary loss, which can be represented as:
@@
where $N$ denotes the total number of training samples, $\hat{s}_{i}$ represents the confidence score of the model’s predicted label $\hat{y}_{i}$ for sample $i$, and $c_{i}=1$ if $\hat{s}_{i}$ equals the true label $y_{i}$; otherwise, $c_{i}=0$. On the other hand, Hebbalaguppe et al. [2022b] provide a more detailed analysis of (6) and propose a Multi-class DCA (MDCA), which offers a more granular examination by considering all classes instead of focusing solely on the predicted label. The proposed auxiliary loss can be written as:
@@
Advantages and challenges. The differentiability of these auxiliary loss terms ((6) and (7)) facilitates their integration with other application-specific loss functions, allowing them to enhance calibration without significantly impacting the primary classification loss, $\ell$. This minimizes any adverse effects on performance while correcting calibration errors. For instance, FL algorithms like FedProx Sahu et al. [2018] and FedNova Wang et al. [2020b] incorporate custom loss functions within their local modeling function, ClientOPT (outlined in Algorithm 1), to address deviations from the global model due to non-IID data. Despite these enhancements, it is critical to note that these advanced FL algorithms still fundamentally require a standard classification loss, which they fine-tune for specific objectives. When integrating calibration into the ClientOPT of FL algorithms, we will exclusively calibrate the standard classification loss, maintaining the integrity of other design elements, such as global aggregation, to ensure their original objectives are not compromised. Choosing an appropriate $\beta_{m}$ tailored to each client $m$ is the remaining challenge.

### v3

@@ 1 Introduction
Federated learning (FL) has rapidly gained traction as a prominent distributed machine learning approach, enabling collaborative model training across a network of clients by periodically aggregating their local models at a central server (Konecný et al., 2016; McMahan et al., 2017). Its significance extends to various critical applications, including medical diagnostics, self-driving cars, and multilingual systems, where data privacy and decentralized training are paramount. FL has been widely studied with a focus on improving local model updates (Reddi et al., 2021; Sahu et al., 2018), refining global aggregation methods (Ji et al., 2019; Wang et al., 2020a), and enhancing communication efficiency (Sattler et al., 2019; Diao et al., 2020). Most of these works in FL consider accuracy as the main performance metric.
@@
Motivation. However, beyond accuracy, in various decision-making scenarios where an incorrect prediction may result in high risk (e.g., medical applications or autonomous driving), it is also crucial for the users to determine whether to rely on the FL model’s prediction or not for each decision. In particular, users should rely on the neural network’s decision only when the prediction is likely to be correct. Otherwise, they may need to consider alternatives such as human decision. To achieve this, the trained FL model should have a reliable confidence in each of its predictions, meaning that the confidence of the neural network matches well with its actual accuracy. In centralized settings, recent studies including Guo et al. (2017) have uncovered that neural networks are often miscalibrated, indicating that the prediction confidence of the model does not accurately reflect the probability of correctness. Handling this miscalibration issue is even more important in many FL use-cases, where an overconfident global model could lead to misinformed decisions with potentially severe consequences for each client.
@@
In centralized training settings, research generally follows two paths to address this miscalibration issue. First, train-time calibration methods (Hebbalaguppe et al., 2022b; Liang et al., 2020; Liu et al., 2021; Mukhoti et al., 2020; Kumar et al., 2018; Müller et al., 2019; Lin et al., 2017) incorporate explicit regularizers during the training process to adjust neural networks, scaling back over/under-confident predictions. On the other hand, post-hoc calibration (Guo et al., 2017; Hendrycks & Gimpel, 2016; Hinton et al., 2015; Kull et al., 2017) transforms the network’s output vector to align the confidence of the predicted label with the actual likelihood of that label for the sample. Post-hoc calibration methods are applied to the already trained model to improve calibration using an additional holdout dataset. However, an appropriate holdout dataset may not be available in many privacy-sensitive applications, making post-hoc calibration strategies often impractical for FL.
@@
Figure 1: Reliability diagrams and calibration errors for centralized training and non-IID FL (using FedAvg) trained with various calibration methods on CIFAR-100 dataset. Our method ensures well-calibrated FL, evidenced by a notably smaller calibration error and a smaller gap (red region) between confidence and accuracy.
@@
Goals and observations. In this paper, we aim to incorporate model calibration into the FL training process, a problem that has been largely underexplored in existing research. This setting introduces new challenges to improving model calibration, as the heterogeneous data distributions across different clients must be carefully considered. We first observe whether existing train-time calibration methods applied during client-side local training can effectively address calibration needs in FL. To gain insights, in Fig. 1, we report accuracy, confidence, and the expected calibration error (ECE) of different methods in different settings. Figs. 1(a) to  1(b) illustrate the capability of train-time calibration methods to narrow the gap between model confidence and accuracy in centralized learning. A comparison between Figs. 1(a) and  1(c) reveals that FL experiences more significant model miscalibration than centralized learning. This disparity is likely due to data heterogeneity across distributed FL clients, causing each client’s local data to have a different impact on calibration performance. When applying auxiliary-based calibration methods, such as DCA (Liang et al., 2020) and MDCA (Hebbalaguppe et al., 2022b) – which add a penalty term to the original classification loss – to FL’s local training (Fig. 1(d)), a better ECE is achieved compared to FL without calibration (Fig. 1(c)). However, we observe that the model in Fig. 1(d) is still not well-calibrated by neglecting global calibration needs in heterogeneous FL settings, resulting in overconfident predictions. Given the fact that train-time calibration methods directly modify confidence during training, potentially affecting the model’s accuracy—which is also a critical factor—we thus pose the following question:
@@
Contributions. In this paper, we propose Non-Uniform Calibration for Federated Learning (NUCFL), a versatile and generic model calibration framework for FL that integrates seamlessly with any existing FL algorithms, such as FedProx (Sahu et al., 2018), Scaffold (Karimireddy et al., 2019), FedDyn (Acar et al., 2021), and FedNova (Wang et al., 2020b). Designed to work with any train-time calibration auxiliary loss, NUCFL is a straightforward yet effective solution that dynamically adjusts calibration penalties based on the relationship between global and local distribution. Specifically, our intuition is that if each client’s local model closely resembles the global model, it is likely that the local model and data represent the global calibration needs well. Therefore, our idea is to choose a large penalty term to improve model calibration for the clients that have a high similarity, while imposing a small penalty for other clients to prioritize accuracy. This non-uniform calibration approach considering data heterogeneity across FL clients significantly minimizes calibration errors without accuracy degradation, as seen in Figure 1(e). Importantly, the flexibility of NUCFL allows for continuous improvement as new FL algorithms or calibration losses are developed. Our key contributions are summarized as follows:
  • Our work pioneers a systematic study on model calibration in FL. We analyze whether existing calibration methods can be applied to FL algorithms and study the following question: How does model calibration impact the behavior of federated optimization methods?
@@
  • Extensive experiments demonstrate that NUCFL is highly flexible, seamlessly integrating with 5 different FL algorithms (e.g., FedProx, Scaffold, FedDyn) and various auxiliary-based train-time calibration methods to improve both calibration performance and accuracy.
@@ 2 Related Work
Federated learning. FedAvg (McMahan et al., 2017) is the pioneering FL algorithm, with numerous adaptations proposed to mitigate the accuracy drop experienced with non-IID data. Much of the research has focused on designing better global aggregation methods to enhance model convergence and performance. Some researchers (He et al., 2020; Lin et al., 2020) replaced weight averaging with model ensemble and distillation, others (Yurochkin et al., 2019; Wang et al., 2020a) matched local model weights before averaging, and another group (Ji et al., 2019) introduced an attention mechanism to optimize global aggregation. Another significant area has concentrated on optimizing local training protocols to optimize computation and data usage on clients’ devices. Some researchers (Sahu et al., 2018; Acar et al., 2021) employed regularization toward the global model, while Karimireddy et al. (2019) utilized server statistics to correct local gradients. Additionally, various studies have addressed challenges specific to FL, such as ensuring fairness across all participants (Li et al., 2020), enhancing communication efficiency to minimize bandwidth usage (Ko et al., 2023), and addressing privacy concerns inherent in distributed computations (Truong et al., 2020). However, the main focus of existing FL research is on improving accuracy. The mismatch between confidence and actual accuracy in these works leads to overconfident FL models, making FL inapplicable in high-risk decision-making scenarios. Our work bridges this gap by investigating an overlooked aspect of federated learning: model calibration.
@@
Model calibration. Model calibration techniques can be categorized into post-hoc calibration and train-time calibration methods. Post-hoc methods adjust the model using a hold-out dataset after training. Common examples include Temperature Scaling (Guo et al., 2017), which adjusts logits by a scalar learned from a hold-out set before softmax, and Dirichlet Calibration (Kull et al., 2017), which modifies log-transformed class probabilities with an extra network layer. However, in FL applications, the lack of an additional validation set due to privacy concerns often renders post-hoc calibration methods inapplicable. On the other hand, train-time calibration integrates calibration directly within the training process. Müller et al. (2019) enhance calibration using label smoothing on soft targets, while  Lin et al. (2017) use focal loss to minimize the KL divergence between predicted and target distributions. Recently, augmenting standard cross-entropy loss with additional penalty loss terms has become a popular train-time calibration method (Liang et al., 2020; Hebbalaguppe et al., 2022b; Kumar et al., 2018). This auxiliary loss penalizes the model when there is a reduction in cross-entropy loss without a corresponding increase in accuracy, often indicative of overfitting. However, as seen in Fig. 1, a direct application of these methods still faces limitations, as they are not able to meet the calibration needs considering data heterogeneity across FL clients. Our work fills this gap by taking a non-uniform calibration approach considering the characteristics of FL.
@@
Calibration and FL. Peng et al. (2024) proposed FedCal, a calibration scaler aggregated solely from local clients, and, to the best of our knowledge, is the only prior work explicitly addressing calibration error in FL. However, FedCal does not consider the interaction between global and local calibration needs, which can lead to a calibration bias towards local heterogeneity. Our calibration framework addresses this issue by considering the interaction between client and server, thereby better accommodating broader global calibration requirements. We empirically confirm these advantages of our approach compared to the existing work in Section 5.2. It is also worth mentioning that Zhang et al. (2022); Luo et al. (2021) also use the term “calibration” in the FL setup. However, we stress that the context of calibration in their work is different from the one in the model calibration literature that considers confidence of the prediction. Zhang et al. (2022); Luo et al. (2021) focuses on calibrating the classifier to improve the accuracy itself, which can be categorized into conventional FL works that do not consider the error between model confidence and accuracy.
@@ 3.1 Federated Learning
In FL, the training data is gathered from a set of clients $M$ and each client $m$ possesses a training set $\mathcal{D}_{m}=\{x_{i},y_{i}\}^{|\mathcal{D}_{m}|}_{i=1}$, where $x$ represents the input data and $y$ denotes the corresponding true label. We consider the following standard optimization formulation of federated training and seek to find model parameters $w$ that solve the problem:
@@
where $w$ represents the model parameter, $\mathcal{D}=\cup_{m\in M}\mathcal{D}_{m}$ denotes the aggregated training set from all clients, and $\mathcal{L}_{m}$ measures the average loss of a model on the training data of the $m$-th client. The objective is to find a model that fits all clients’ data well on (weighted) average.
FL methods involve parallel local training at clients and global aggregation at a server over multiple communication rounds to address the aforementioned problem. Most federated optimization methods fit within a general framework outlined by Reddi et al. (2020), as shown in Algorithm 1. At round $t$, the server sends its previous global model ${w}^{(t-1)}$ to all clients as initialization. Each client $m$ then performs $E$ epochs of local training using ClientOPT depending on the FL algorithm, and subsequently produces a local model ${w}_{m}^{(t,E)}$. Then each client communicates the difference $\delta^{(t)}_{m}$ between their learned local model and the server model, denoted as $\delta^{(t)}_{m}={w}^{(t-1)}-{w}^{(t,E)}_{m}$. The server computes a weighted average of the client updates, denoted as $\delta^{(t)}$, and then updates the global model ${w}^{(t)}$ using ServerOPT, which also depends on FL designs. For example, FedAvg employs standard stochastic gradient descent updates for ClientOPT, and its ServerOPT is formulated as $w^{(t)}=w^{(t-1)}-\delta^{(t)}$.
@@ 3.2 Model Calibration
In supervised multi-class classification, the input $x\in\mathcal{X}$ and label $y\in\mathcal{Y}=\{1,...,K\}$ are random variables following a ground truth distribution $\pi(x,y)=\pi(y|x)\pi(x)$. Let $s$ be the confidence score vector in $\mathbb{R}^{K}$, with $s[y]=f(x)$ representing the confidence that the model $f(\cdot)$ predicts for a class $y$ given input $x$. The predicted class $\hat{y}$ is computed as: $\hat{y}=\arg\max_{y}s[y]$, with the corresponding confidence $\hat{s}=\max_{y}s[y]$. A model is considered perfectly calibrated (Guo et al., 2017) if:
@@
Note that “increasing confidence” refers to adjusting the model’s predicted probabilities to higher values, but it does not necessarily imply “better model calibration.” Excessively high confidence can lead to overconfident predictions, potentially misleading users to over-rely on the model even when its predictions are inaccurate. The primary goal of model calibration, whether via train-time or post-hoc methods, is to minimize discrepancy (3) to ensure the model’s confidence accurately reflects its prediction accuracy. This ensures that users can directly know the probability of correctness based on the model confidence, allowing the model to be applied in various decision-making scenarios.
@@ 4.1 Bridging Federated Learning and Model Calibration
where $w_{m}$ is the local model, $t$ the current global round, $e$ the local iteration index, and $\ell(\cdot)$ the loss function (e.g., cross-entropy) applied to a data instance. One can also modify the loss in (4), i.e., by adding a proximal term as in FedProx (Sahu et al., 2018). Recent studies on calibration (Liang et al., 2020; Hebbalaguppe et al., 2022b) suggest that incorporating an auxiliary loss, which captures the disparity between accuracy and confidence (as in equation (3)), and penalizing the model accordingly, leads to superior performance compared to direct loss type adjustments (Müller et al., 2019; Brier, 1950; Lin et al., 2017). These auxiliary loss methods maintain the original classification loss and seamlessly add a penalty, minimizing interference with FL performance, especially since some FL algorithms are specifically designed around their classification loss strategies. This makes the auxiliary loss methods easily applicable and effective across various FL settings.
@@
where $\ell_{cal}(\cdot)$ is the auxiliary calibration loss function, and $\beta_{m}$ is a scalar weight for each client $m$. In the centralized calibration case, this scalar weight is manually selected to refine the auxiliary calibration loss. However, in our design, we automatically determine $\beta_{m}$ based on the relationship between the server and clients in a FL setup (without any hyperparameter tuning), taking into account the heterogeneity as detailed in Section 4.2. The auxiliary calibration loss $\ell_{cal}$ can be derived from various calibration designs. For instance, Liang et al. (2020) directly adds the difference between confidence and accuracy (DCA) as an auxiliary loss, which can be represented as:
@@
where $N$ denotes the total number of training samples, $\hat{s}_{i}$ represents the confidence score of the model’s predicted label $\hat{y}_{i}$ for sample $i$, and $c_{i}=1$ if $\hat{s}_{i}$ equals the true label $y_{i}$; otherwise, $c_{i}=0$. On the other hand, Hebbalaguppe et al. (2022b) provide a more detailed analysis of (6) and propose a Multi-class DCA (MDCA), which offers a more granular examination by considering all classes instead of focusing solely on the predicted label. The proposed auxiliary loss can be written as:
@@
Advantages and challenges. The differentiability of these auxiliary loss terms ((6) and (7)) facilitates their integration with other application-specific loss functions, allowing them to enhance calibration without significantly impacting the primary classification loss, $\ell$. This minimizes any adverse effects on performance while correcting calibration errors. For instance, FL algorithms like FedProx (Sahu et al., 2018) and FedNova (Wang et al., 2020b) incorporate custom loss functions within their local modeling function, ClientOPT (outlined in Algorithm 1), to address deviations from the global model due to non-IID data. Despite these enhancements, it is critical to note that these advanced FL algorithms still fundamentally require a standard classification loss, which they fine-tune for specific objectives. When integrating calibration into the ClientOPT of FL algorithms, we will exclusively calibrate the standard classification loss, maintaining the integrity of other design elements, such as global aggregation, to ensure their original objectives are not compromised. Choosing an appropriate $\beta_{m}$ tailored to each client $m$ is the remaining challenge.